Suppose that left-handedness is independent of possession of a Visa card. Suppose that the probability that a person is left-handed is .10 and that the probability a person possesses a Visa card is .60. What is the probability that a person possesses a Visa card or that they are left-handed?
Answer #1

Let , A be the event that the person is left-handed and B be the event that the person possesses a Visa card.

Given that P(A)=0.10 and P(B)=0.60

Since , the event A and B are independent.

By definition of independence of event , P(ANB) = P(A) * P(B)

Now we want to find the probability that a person possesses a Visa card or that they are left-handed.

By using the addition rule ,

P(AUB) = P(A) + P(B) - P(ANB)

= P(A) + P(B) - P(A)*P(B)

=0.10 +0.60 -0.10 +0.60

=0.64